Can you please help me with this problem? I am a little slow so you will need to break it down step by step. Thank you so much.
The square root of x^2 to the third power =9

Answer by stanbon(48568)

The square root of x^2 to the third power =9
--------
[(sqrt(x^2)]^3=9
x^3=9
x=9^(1/3)
Cheers,
Stan H.